# A Synergistic Framework for Learning Shape Estimation and Shape-Aware Whole-Body Control Policy for Continuum Robots

### v3

@@ Abstract
In this paper, we present a novel synergistic framework for learning shape estimation and a shape-aware whole-body control policy for tendon driven continuum robots. Our approach leverages the interaction between two Augmented Neural Ordinary Differential Equations (ANODEs) — the Shape-NODE and Control-NODE — to achieve continuous shape estimation and shape-aware control. The Shape-NODE integrates prior knowledge from Cosserat rod theory, allowing it to adapt and account for model mismatches, while the Control-NODE uses this shape information to optimize a whole-body control policy, trained in a Model Predictive Control (MPC) fashion. This unified framework effectively overcomes limitations of existing data-driven methods, such as poor shape awareness and challenges in capturing complex nonlinear dynamics. Extensive evaluations in both simulation and real-world environments demonstrate the framework’s robust performance in shape estimation, trajectory tracking, and obstacle avoidance. The proposed method consistently outperforms state-of-the-art end-to-end, Neural-ODE, and Recurrent Neural Network (RNN) models, particularly in terms of tracking accuracy and generalization capabilities. The code and pretrained models are available at https://github.com/MohammadKasaei/WholeBodyControl_CTR.
@@ I Introduction
Soft continuum robots are constructed from flexible materials such as rubber, silicone, or elastomers, allowing them to conform to surfaces and objects while maintaining a level of durability that rigid robots do not possess [1, 2]. This adaptability makes them well-suited for diverse applications, including medical procedures [3], search and rescue operations [4], and exploration missions [5]. Despite these advantages, designing and controlling soft continuum robots remain challenging [6, 7]. These challenges stem from the difficulty of predicting their shapes due to nonlinear behaviors and the need to model complex structural deformations with high degrees of freedom, making precise whole-body control a particularly complex task [8].
@@ I-A Shape Estimation
The shape of continuum robots is typically described by kinematic or dynamic models that relate the configuration of the robot’s backbone to its joint inputs. Various mathematical approaches have been developed for modeling continuum robots, including Piecewise Constant Curvature (PCC) [9, 10], polynomial curvature fitting [11], lumped parameter models [12, 13], and reduced-order finite element models [14, 15]. A detailed review of these methodologies is provided in [16]. One of the most widely used frameworks for modeling continuum robots is Cosserat rod theory [17, 18, 19, 20, 21]. However, shape estimation based on model-driven approaches often suffers from significant inaccuracies due to the dependence on external force information, which is often unavailable [22]. Additionally, the inherent challenges posed by highly non-linear kinematics, torsion, and friction further compromise the accuracy of these models [23]. Furthermore, these methods can be computationally intensive and may not adequately represent the complex nonlinear dynamics inherent to soft robots.
@@ IV-B Experiments
Two experimental scenarios were designed to evaluate the performance of the proposed framework on the real robot. The first scenario focuses on trajectory tracking, where the robot is tasked with following various trajectories, consistent with those used in the simulations. The second scenario is designed to assess the robustness of the method by requiring the robot to follow a helical trajectory while carrying additional payloads, with weights ranging from 5 to 20 grams (see Fig. 6). A video demonstrating the experiments and results is available online at https://youtu.be/SIR7hVi_Bd0.
@@ VI Conclusion
This paper introduced a synergistic framework that combines shape estimation and shape-aware control for continuum robots using two Augmented Neural Ordinary Differential Equations (ANODEs). The integration of the Shape-NODE, with its built-in model of Cosserat rod theory, allowed for accurate shape prediction and adaptation to model mismatches, while the Control-NODE enabled the development of a shape-aware whole-body control policy optimized in an MPC fashion. Our extensive simulations and experiments validate the effectiveness and robustness of the framework, especially in complex scenarios involving trajectory tracking and obstacle avoidance. The proposed framework consistently outperformed alternative approaches, like end-to-end models, Neural ODEs, and RNNs, both in terms of tracking accuracy and generalization to unseen data.

### v4

@@ Abstract
In this paper, we present a novel synergistic framework for learning shape estimation and a shape-aware whole-body control policy for tendon driven continuum robots. Our approach leverages the interaction between two Augmented Neural Ordinary Differential Equations (ANODEs) — the Shape-NODE and Control-NODE — to achieve continuous shape estimation and shape-aware control. The Shape-NODE integrates prior knowledge from Cosserat rod theory, allowing it to adapt and account for model mismatches, while the Control-NODE uses this shape information to optimize a whole-body control policy, trained in a Model Predictive Control (MPC) fashion. This unified framework effectively overcomes limitations of existing data-driven methods, such as poor shape awareness and challenges in capturing complex nonlinear dynamics. Extensive evaluations in both simulation and real-world environments demonstrate the framework’s robust performance in shape estimation, trajectory tracking, and obstacle avoidance. The proposed method consistently outperforms state-of-the-art end-to-end, Neural-ODE, and Recurrent Neural Network (RNN) models, particularly in terms of tracking accuracy and generalization capabilities. The code and pretrained models are available at https://github.com/SIRGLab/WholeBodyControl_CTR.
@@ I Introduction
Soft continuum robots are constructed from flexible materials such as rubber, silicone, or elastomers, allowing them to conform to surfaces and objects while maintaining a level of durability that rigid robots do not possess [1, 2]. This adaptability makes them well-suited for diverse applications, including medical procedures [3, 4, 5], search and rescue operations [6], and exploration missions [7]. Despite these advantages, designing and controlling soft continuum robots remain challenging [8, 9]. These challenges stem from the difficulty of predicting their shapes due to nonlinear behaviors and the need to model complex structural deformations with high degrees of freedom, making precise whole-body control a particularly complex task [10].
@@ I-A Shape Estimation
The shape of continuum robots is typically described by kinematic or dynamic models that relate the configuration of the robot’s backbone to its joint inputs. Various mathematical approaches have been developed for modeling continuum robots, including Piecewise Constant Curvature (PCC) [11, 12], polynomial curvature fitting [13], lumped parameter models [14, 15], and reduced-order finite element models [16, 17]. A detailed review of these methodologies is provided in [18]. One of the most widely used frameworks for modeling continuum robots is Cosserat rod theory [19, 20, 21]. However, shape estimation based on model-driven approaches often suffers from significant inaccuracies due to the dependence on external force information, which is often unavailable [22]. Additionally, the inherent challenges posed by highly non-linear kinematics, torsion, and friction further compromise the accuracy of these models [23]. Furthermore, these methods can be computationally intensive and may not adequately represent the complex nonlinear dynamics inherent to soft robots.
@@ IV-B Experiments
Two experimental scenarios were designed to evaluate the performance of the proposed framework on the real robot. The first scenario focuses on trajectory tracking, where the robot is tasked with following various trajectories, consistent with those used in the simulations. The second scenario is designed to assess the robustness of the method by requiring the robot to follow a helical trajectory while carrying additional payloads, with weights ranging from 5 to 20 grams (see Fig. 6). A video demonstrating the experiments and results is available online. [^1] https://youtu.be/aZWfDQK9hpI?si=52BSPRC4p7HSqjsl
@@ VI Conclusion
This paper introduced a synergistic framework that combines shape estimation and shape-aware control for continuum robots using two Augmented Neural Ordinary Differential Equations (ANODEs). The integration of the Shape-NODE, with its built-in model of Cosserat rod theory, allowed for accurate shape prediction and adaptation to model mismatches, while the Control-NODE enabled the development of a shape-aware whole-body control policy optimized in an MPC fashion. Our simulations and experiments validate the effectiveness and robustness of the framework, especially in complex scenarios involving trajectory tracking and obstacle avoidance. The proposed framework consistently outperformed alternative approaches, like end-to-end models, Neural ODEs, and RNNs, both in terms of tracking accuracy and generalization to unseen data.